OpenStudy (anonymous):

What is the value of... 4*integral [ sqrt ( 1 - x^2) ]dx from 0 to 1

OpenStudy (experimentx):

this is just area of circle with unit radius.

OpenStudy (experimentx):

this expression gives you the area of circle with radius 'a'. Make use of trigonometric substitution \(x = a \cos \theta \) to get the result. \[ 4\times \int_0^a \sqrt{a^2 - x^2}dx\]

OpenStudy (anonymous):

So the expression's value is pi, correct?

OpenStudy (anonymous):

How would you solve the integral without the geometry?

OpenStudy (anonymous):

From what I remember of trig: 1 - cos^2 (x) = sin^2 (x), so sqrt [ sin^2 ] = sinx Therefore, we can simplify the expression into 4*integral [ sinx ]dx from 0 to 1 int [ sinx ] = -cosx From this point, QED